7 of 10 A 8 85 g sample of hydrate of sodium carbonate, Na,CO, XH,0 was heated caretuly until no more mass was lost from the sample. Ater heating, the final weight of the material was 7.55 g What is the sodium carbonate (Na CO,)-to-water mole ratio? O1:1 01:0 〇2:1 012
Answer #1

Given

weight of hydrated material = 8.55 g

Weight after losing hydrate = 7.55 g (only Na2CO3)

Molar mass of Na2CO3 = 106 g/mol

No. of moles of Na2CO3 = 7.55 g / 106 g/mol = 0.0712 moles

same no. of moles of Na2CO3 with xH2O will be present in intial material

so Molar mass of initial material = 8.55 g / 0.0712 moles = 120 g/mol

Molar mass of Na2CO3.xH2O = 120 g/mol

Molar mass of xH2O = 120 - 106 = 18 g/mol

which is molar mass of 1 H2O so

molar ratio is 1:1 answer 1:1